A factory with capacity of 700,000 units per year operates at 62% capacity. The annual income is $430,000.00. Annual fixed costs are $190,000.00 and the variable costs are $0.348 per unit.

a) What is the current loss or profit?

b) What is the breakeven point?
Answer #1

Answer

a)

Profit = Total Revenue - Total Cost

$430,000

Total Cost = fixed Cost + Variable Cost

Variable Cost = 0.348*Q and Q = 62% of 700,000 = 0.62*700,000 = 434,000

=> 0.348*434000 = $151032

Total Cost = 190,000 + 151032 = 341032

Hence Profit = 430,000 - 341032 = 88968

b)

AS they sold 43400 and earned income = 430,000.

Hence Per unit price = 430,000/434000 = 0.991

In order to be breakeven

Total Cost = Total Revenue

=>0.991Q = 0.348Q + 190000

=> Q = 295490

Hence In order to be breakeven they should produce 295490 units